What is the ∆S of N2(g) + 3H2(g) ---> 2NH3(g)?


Given


N2(g)

(J/mol x K) = 191.5


H2(g)

(J/mol x K) = 130.6


NH3(g)

(J/mol x K) = 192.3


1
Expert's answer
2022-04-26T13:19:04-0400

At 298 K as a standard temperature:

S°(N2, g) = 191.5 J mol−1 K−1

S°(H2, g) = 130.6 J mol−1 K−1

S°(NH3, g) = 192.3 J mol−1 K−1


Solution:

The entropy change (ΔS°) in a chemical reaction is given by the sum of the entropies of the products minus the sum of the entropies of the reactants.

ΔS° =∑nS°(products) − ∑mS°(reactants)


Balanced chemical equation:

N2(g) + 3H2(g) → 2NH3(g)


From the balanced equation we can write the equation for ΔS° (the change in the standard molar entropy for the reaction):

ΔS° = 2×S°(NH3, g) − [S°(N2, g) + 3×S°(H2, g)]

ΔS° = 2×192.3 − [191.5 + 3×130.6] = −198.7 (J mol−1 K−1)

ΔS° = −198.7 J mol−1 K−1


Answer: The entropy change is −198.7 J mol−1 K−1